OpenStudy (anonymous):

A common inhabitant of human intestines is the bacterium Escherichia coli. A cell of this bacterium in a nutrient-broth medium divides into two cells every 20 minutes. The initial population of a culture is 73 cells.

OpenStudy (anonymous):

Welcome to OpenStudy man!

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

So what is the question? How many are there after a certain amount of time?

OpenStudy (anonymous):

P(t)=p(0)e^kt then P(1)=73e^8t. therefore k = ln(8)

OpenStudy (anonymous):

is that right and can you explain the logic behind it

OpenStudy (anonymous):

and then how do i calulate the rate of change after 8 hours

OpenStudy (anonymous):

Plug in 8 for t, and solve for P(t). I can explain better in a moment.

OpenStudy (anonymous):

k i'll wait for you to come back i thought 8 was getting plugged in for k. time would be 20 minutes i thought.